The manufacturer of cans of salmon that are supposed to have a net weight of 6 ounces tells you that the net weight is actually a normal random variable with a mean of 6.1 ounces and a standard deviation of 0.22 ounce. Suppose that you draw a random sample of 26 cans. Find the probability that the mean weight of the sample is less than 6.09 ounces.

Probability =
Answer #1

Solution :

Given that ,

mean = = 6.1

standard deviation = = 0.22

n = 26

= 6.1

=  / n = 0.22 / 26=0.04

P( <6.09 ) = P[( - ) / < (6.09-6.1) / 0.04]

= P(z < -0.25)

Using z table  

Probability == 0.4013